# Exploring the effect of different cosmologies on the Epoch of Reionization 21-cm signal with Polar

### v1

@@ 3.1.1 Galactic properties
In Figure 5, we present the mass-binned star formation rate (SFR) of galaxies in the four models at $z=10$ and 9, with the top panel showing the unconstrained case. The solid lines are the median SFR, with the shaded regions referring to the 16th to 84th percentiles. The star formation main sequence (SFMS) in each case is additionally compared with results from various recent JWST programs represented by grey circles with a binning of ±$\pm 0.25$ around the redshift. As all parameters affecting the star formation rate were kept the same across all four models, we note that the overall trends are similar, and agree with observations. The impact of matter clustering is seen at the low-mass end, where a larger clustering allows for the formation of a statistically significant sample of smaller mass galaxies in the $\sigma_{8}$ high and $h$ high models, while the smallest mass galaxies formed in the $\sigma_{8}$ low case are an order of magnitude more massive. Further, at the high mass end, we note that the $\sigma_{8}$ high model produces the most massive galaxies, which also have the highest SFR.
@@
Figure 6: A and B-series averaged star formation efficiency (SFE) as a function of halo mass for the unconstrained (top) and constrained (bottom) cases, for the fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green) models at $z=10$ and 9. Solid lines refer to the median SFE in each mass bin, with 16th and 84th percentiles shown as shaded regions. Grey circles are Spitzer observations from Stefanon et al. (2021), and grey triangles are abundance matching estimates from Tacchella et al. (2018).
To confirm the agreement in star formation rates, in Figure 6 we additionally look at the global star formation efficiency (SFE) at the same redshifts with the top panel showing the unconstrained case. We define SFE as the ratio of the stellar and halo mass scaled by $f_{\rm b}=\Omega_{\rm b}/\Omega_{\rm m}$, i.e. ${\rm M}_{\rm star}/{\rm M}_{\rm halo}f_{\rm b}$, as a function of the halo mass ${\rm M}_{\rm halo}$. We also compare this to Spitzer observations at $z=7$ (Stefanon et al., 2021) and to abundance matching estimates from Tacchella et al. (2018). As mentioned above, the parameters controlling the star formation efficiency are the same across all models, and thus the corresponding global SFE is the same as well. The only difference is in the largest galaxies formed, which is directly proportional to $\sigma_{8}$ as it controls the clustering of mass. We note that the SFE is marginally higher than the abundance matching estimates for the lower mass halos, but the difference is $\lesssim 0.3$ dex. Further, the weak mass dependence of the global SFE trend is in agreement with zoom-in hydrodynamical simulations like FIREboxHR (Feldmann et al., 2024).
@@
We note that despite the disagreement in UVLFs for the four cosmological models shown in the top panel of Figure 2, the models still reproduce fairly well observable quantities such as the SFR and global SFE, as shown in Section 3.1.1. Further, they are broadly in agreement with available data. We now look at IGM observables, and more specifically at the 21-cm signal power spectrum (calculated according to Equation 4), shown in the top panel of Figure 7 for the unconstrained case at $z=9$. At this redshift, we note from the top panel of Figure 4 that the reionization history is primarily governed by the matter clustering. Thus, the $\sigma_{8}$ low model having the highest $\langle x_{\rm HI}\rangle$ suggests a significant contrast between its large neutral and ionized regions, thus leading to a higher power amplitude up to $k=0.40~{}h\rm cMpc^{-1}$ as compared to the other three models. On the other hand, the greater matter clustering in the $\sigma_{8}$ high and $h$ high models leads to lower contrast at these scales due to more numerous ionized regions, and thus less power. However, at smaller scales, having more ionized regions allows for greater contrast, and thus the $h$ high model catches up with the fiducial model, and the $\sigma_{8}$ high model has the highest power overall at $k>0.40~{}h\rm cMpc^{-1}$.
@@ 3.2.1 Galactic properties
To further explore the impact of the energy released by supernovae, in the bottom panel of Figure 6 we also look at the global SFE for the constrained case. Here, we notice that, despite similar SFRs at the high-mass end, the efficiency of star formation in each case differs. The SFE for the $\sigma_{8}$ low model is significantly boosted compared to observations, while the fiducial and $h$ high models only mildly overpredict the observed SFE. The $\sigma_{8}$ high case, on the other hand, is the best match. We note that the global SFE is inversely proportional to $E_{\rm SN}$ for all halo masses. The weak mass dependence trend of the SFE is the same of the unconstrained case, and in agreement with the trends of the FIREboxHR simulations (Feldmann et al., 2024) for the fiducial, $\sigma_{8}$ high and $h$ high models, while the $\sigma_{8}$ low model is an order of magnitude higher but with the same slope.
@@
While the UVLFs at $M_{\rm 1600,AB}<-18$ in the constrained case are similar for all four cosmological models, the SFR and global SFE show differences, and in some cases do not match observations. In such a scenario, it becomes interesting to analyze the impact on the 21-cm signal, which is presented in the bottom panel of Figure 7 at $z=9$ for the constrained case. Unlike the unconstrained case where the differences in matter clustering contributed to differences not only in the overdensity term but also in the neutral hydrogen fraction term in equation 2, here the changes in astrophysical parameters lead to a drastic reduction in the differences between the value of $\langle x_{\rm HI}\rangle$ across the four models. Thus, the differences in the 21-cm signal power spectra of the four models are governed primarily by the overdensity term. This leads to a higher power across all wave modes for the $\sigma_{8}$ high model due to greater matter clustering. While one would expect the $h$ high model to have slightly higher power than the fiducial one due to the dependence on overdensity, the neutral-fraction term in Equation 2 reduces the contrast due to a marginally faster reionization process, so that the power in the two cases becomes similar. Even the $\sigma_{8}$ low model matches the fiducial and $h$ high models’ power spectra due to a faster reionization process.
In the bottom panel of Figure 8, we present the redshift evolution of the 21-cm signal power spectrum at $k=0.15~{}h\rm cMpc^{-1}$ for the constrained case and notice a significant reduction in the differences of $\Delta^{2}_{\rm 21cm}(k=0.15~{}h\rm cMpc^{-1})$ values between the various cosmological models at $512>z>$ 5 compared to the extent of differences in the unconstrained case (top panel of Figure 8). Indeed, at $8.510>z>$ 8.5, where LOFAR observations are focused, variations in the $\sigma_{8}$ parameter lead to differences in the power spectrum of only a few mK2. While the exact magnitude of the difference is subject to the position of the peak of $\Delta^{2}_{\rm 21cm}(k=0.15~{}h\rm cMpc^{-1})$, which in turn is governed by modeling assumptions like the choice of $f_{\rm esc}$ (see Appendix C for a comparison between $f_{\rm esc}=12.5\%$ and 25%), the key takeaway is that significantly different choices of cosmological and astrophysical parameters can still lead to similar 21-cm signal observables, even when constrained by the UVLFs from JWST and HST observations. Thus, to explore the parameter space of models that agree with upper limits of 21-cm signal power spectra observations at the redshifts of interest to LOFAR, one should still consider various cosmological and astrophysical parameters as free. However, jointly constraining these models with a more diverse set of galactic observables from JWST and upcoming surveys with Euclid and SPHEREx can allow us to narrow down the choice of viable ones. Additionally, tomographic imaging of the distribution of neutral Hydrogen with the upcoming SKA-Low could also be used for additional constraints on IGM properties from images such as those shown in Figure 3 (see for e.g. Bianco et al., 2021).
@@ 4 Discussion
By analyzing various galactic and IGM observables, we note that the interplay of astrophysical and cosmological parameters can lead to similarities as well as differences. For example, in the constrained case, while agreement with observed UVLFs leads to similar SFR for the high-mass galaxies across all models, there are significant differences in the SFR and global SFE for low-mass galaxies. On the other hand, while in the unconstrained case the SFR and global SFE show a very similar dependence on stellar mass in all models because the parameters that regulate such quantities are the same, the UVLFs exhibit significant differences.
@@
These similarities in terms of 21-cm observables have a profound impact on inference modeling, as models with very different cosmological and astrophysical parameters may still produce 21-cm signal power spectra in agreement with current upper limits (for example from Mertens et al., 2020; Trott et al., 2020; HERA Collaboration et al., 2023; Acharya et al., 2024c), or possibly also with an eventual measurement. This is because, while the end of reionization may differ among models, at $z>7$ the extent and speed of reionization are still largely similar. This means that fixing a priori any of these parameters could exclude viable models. While introducing cosmological and astrophysical parameters as free parameters during inference modeling has been attempted in several earlier studies (see e.g. Kern et al., 2017; Schneider et al., 2023), these have used only approximate reionization simulation codes due to the significant computational costs involved in building such a high-dimensional parameter space. Indeed, in a more physical approach, it is necessary to run a large number of dark-matter only simulations with different cosmologies, each post-processed with a large set of astrophysical parameters. The problem worsens if one would like to boost the resolution of the simulations to take into account the role of mini halos (see Haiman et al., 2001; Iliev et al., 2005), dwarf galaxies (Wu & Kravtsov, 2024), or Lyman Limit Systems (Georgiev et al., 2024; Giri et al., 2024) which likely play a significant role in the EoR. While faster semi-numerical codes and emulators can be employed, they are not ideal tools to model galaxy-scale physics and radiative transfer effects.
@@ 5 Summary
In this work, we have investigated the impact of different cosmological models on the 21-cm signal and some galactic properties using Polar, which combines $N$-body dark matter simulations run with Gadget-4 with the semi-analytic model of galaxy formation L-Galaxies, and the 1-D radiative transfer code Grizzly. We have applied the framework to four different cosmological models: a “fiducial” model, which adopts values of the cosmological parameters from Planck Collaboration et al. (2020); a “$h$ high” model, with $h=0.733$ based on results from studies of Cepheid variables in the host galaxies of 42 Type Ia supernovae (Riess et al., 2022); a “$\sigma_{8}$ low” case, with $\sigma_{8}=0.702$ from an anisotropic galaxy clustering measurement analysis done by Tröster et al. (2020); and a “$\sigma_{8}$ high” case, with $\sigma_{8}=0.880$ according to recent eROSITA results (Ghirardini et al., 2024). We additionally used the Fixed & Paired approach (Angulo & Pontzen, 2016) to suppress cosmic variance by boosting the effective volumes of the simulations. For all the quantities analyzed, we took an average of the fixed initial conditions case (which we refer to as the A-series) and its corresponding pair (referred to as the B-series).
We then choose astrophysical parameters in L-Galaxies such that the fiducial model matches UV luminosity functions from HST legacy fields and JWST programs at $z=10$ and 9. To investigate the effect of adopting different values for $h$ and $\sigma_{8}$ on other galactic and IGM observables, we use the same astrophysical parameters also for the other three cosmological models. We refer to this as the “unconstrained” case, as by not changing astrophysical parameters for the different cosmological models the resulting UVLFs will not necessarily be consistent with the observed ones. We also built a “constrained” case, where we instead choose the astrophysical parameters such that the UVLFs obtained in each cosmological model are consistent with those observed at $z=10$ and 9. For this, we increased the star formation efficiency and reduced the energy released per supernova in the $\sigma_{8}$ low model to boost star formation and, in turn, its UVLF. On the other hand, we increased the energy released per supernova for the $h$ high and $\sigma_{8}$ high models to suppress star formation and mitigate the impact of increased matter clustering as compared to the fiducial model. For the radiative transfer calculations, we only took into account the stellar sources while modeling the spectral energy distributions and chose a global escape fraction of 12.5% across all redshifts for all four models.
@@
  • Global Star Formation Efficiency (SFE): As the parameters controlling star formation are the same in the unconstrained case, the global SFE at all masses is similar in all models. The SFE is also in agreement with observations and abundance matching estimates. However, in the constrained case, due to a higher value of the parameter regulating the star formation efficiency, the global SFE in the $\sigma_{8}$ low model is also higher. The other three models show minor differences, governed by the differences in the energy released per supernova.
@@
  • 21-cm signal power spectrum: In the unconstrained case, the differences between the four models are not only due to the difference in overdensity but also to those in the neutral fraction and its redshift evolution. However, in the constrained case, the neutral fraction at $z>8$ is similar in all models, as the impact of matter clustering on the neutral fraction is canceled out by the impact of setting different values for the astrophysical parameters controlling star formation. Thus, the power spectrum at higher redshifts is mostly dictated by the overdensity term, and because of this it shows smaller differences among the four models.
@@
Overall, we conclude that different values of cosmological and astrophysical parameters can lead to differences in some observables (e.g. low-mass SFR and global SFE), while others are largely unaffected (e.g. UVLFs and 21-cm signal power spectrum). In particular, we note that despite significantly different galactic processes and reionization histories, the 21-cm power spectra are very similar in power across $k$-bins and in agreement with current observational upper limits (Acharya et al., 2024c; Mertens et al., 2020; Trott et al., 2020; HERA Collaboration et al., 2023). Due to this, when doing inference modeling, it is essential to consider all cosmological and astrophysical parameters as free parameters, with other observational constraints serving as priors. While a limited exploration of astrophysical parameters can be done with excursion set algorithms, a semi-analytic model provides a significantly more rigorous and physical approach to modeling galactic properties. For varying cosmological parameters though, it is necessary to run a large number of $N$-body simulations to populate the prior parameter space. However, running so many high-resolution simulations would be prohibitively expensive in terms of computational resources and time. To address this issue, in future work we intend to incorporate techniques for boosting the resolution of less costly, low-resolution simulations using either analytic or machine learning techniques. Such an implementation would also additionally allow us to build broader training sets of power spectra for the Machine Learning kernels used with Gaussian Process Regression by the LOFAR EoR Key Science Project team (as done by Acharya et al., 2024a).
@@
Lastly, to achieve a stronger constraining power on a broadened parameter space, it is essential to jointly employ a variety of galactic observables, e.g. from JWST, Euclid and upcoming SPHEREx, as well as tomographic imaging of neutral Hydrogen in the IGM with the future SKA-Low.

### v2

@@ 3.1.1 Galactic properties
In Figure 5, we present the mass-binned star formation rate (SFR) of galaxies in the four models at $z=10$ and 9, with the top panel showing the unconstrained case. The solid lines are the median SFR, with the shaded regions referring to the 16th to 84th percentiles. The star formation main sequence (SFMS) in each case is additionally compared with results from various recent JWST programs represented by grey circles with a binning of $\pm 0.25$ around the redshift. As all parameters affecting the star formation rate were kept the same across all four models, we note that the overall trends are similar, and agree with observations. The impact of matter clustering is seen at the low-mass end, where a larger clustering allows for the formation of a statistically significant sample of smaller mass galaxies in the $\sigma_{8}$ high and $h$ high models, while the smallest mass galaxies formed in the $\sigma_{8}$ low case are an order of magnitude more massive. Further, at the high mass end, we note that the $\sigma_{8}$ high model produces the most massive galaxies, which also have the highest SFR.
@@
Figure 6: A and B-series averaged stellar fraction (which serves as a proxy for the star formation efficiency) as a function of halo mass for the unconstrained (top) and constrained (bottom) cases, for the fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green) models at $z=10$ and 9. Solid lines refer to the median stellar fraction in each mass bin, with 16th and 84th percentiles shown as shaded regions. Grey circles are Spitzer observations from Stefanon et al. (2021), and grey triangles are abundance matching estimates from Tacchella et al. (2018).
To confirm the agreement in star formation rates, in Figure 6 we additionally look at the stellar fraction, which serves as a proxy for the global star formation efficiency (SFE) at the same redshifts with the top panel showing the unconstrained case. We define the stellar fraction as the ratio of the stellar and halo mass scaled by $f_{\rm b}=\Omega_{\rm b}/\Omega_{\rm m}$, i.e. ${\rm M}_{\rm star}/{\rm M}_{\rm halo}f_{\rm b}$, as a function of the halo mass ${\rm M}_{\rm halo}$. We also compare this to Spitzer observations at $z=7$ (Stefanon et al., 2021) and to abundance matching estimates from Tacchella et al. (2018). As mentioned above, the parameters controlling the star formation efficiency are the same across all models, and thus the corresponding stellar fraction is the same as well. The only difference is in the largest galaxies formed, which is directly proportional to $\sigma_{8}$ as it controls the clustering of mass. We note that the stellar fraction is marginally higher than the abundance matching estimates for the lower mass halos, but the difference is $\lesssim 0.3$ dex. Further, the weak mass dependence of their trend is in agreement with zoom-in hydrodynamical simulations like FIREboxHR (Feldmann et al., 2024) and ColdSIM (Maio & Viel, 2023).
@@
We note that despite the disagreement in UVLFs for the four cosmological models shown in the top panel of Figure 2, the models still reproduce fairly well observable quantities such as the SFR and the stellar fraction, as shown in Section 3.1.1. Further, they are broadly in agreement with available data. We now look at IGM observables, and more specifically at the 21-cm signal power spectrum (calculated according to Equation 4), shown in the top panel of Figure 7 for the unconstrained case at $z=9$. In particular, we focus on the $0.05<k<0.5h\rm cMpc^{-1}$ regime, which is where the LOFAR telescope is the most sensitive. At redshift $z=9$, we note from the top panel of Figure 4 that the matter clustering primarily governs the reionization history. Thus, the $\sigma_{8}$ low model having the highest $\langle x_{\rm HI}\rangle$ suggests that the power amplitude up to $k=0.40~h\rm cMpc^{-1}$ for it is higher than the other three models, as it is primarily driven by the matter power spectrum. Additionally, the power at these scales is further boosted, because while $\langle x_{\rm HI}\rangle\approx 1$, there are still voxels where it is $\ll 1$ as they are included in ionized regions. These ionized regions in the $\sigma_{8}$ low model are small and few enough to lead to higher contrast on average in the differential brightness temperature across all scales greater than the size of the ionized bubbles. In other words, as matter clustering translates into the neutral hydrogen fraction term of Equation 2 as well as the overdensity term, in a scenario where the ionized regions are not large and numerous enough to wash out contrast, it leads to a marginal boost in power.
The lower $\langle x_{\rm HI}\rangle$ in the other three models suggests suppression of the power spectrum for the 21-cm signal as compared to the matter power spectrum due to more numerous and larger ionized regions, especially in the $\sigma_{8}$ high and $h$ high models. This is due to the washing out of the contrast driven purely by the large-scale structure. However, the more numerous ionized regions present at smaller scales induce on average a greater contrast, and thus the $h$ high model reaches values similar to those of the fiducial model, and the $\sigma_{8}$ high model has the highest power overall at $k>0.40~h\rm cMpc^{-1}$.
@@ 3.2.1 Galactic properties
To further explore the impact of the energy released by supernovae, in the bottom panel of Figure 6, we also look at the global SFE for the constrained case by again considering the stellar fraction. Here, we notice that, despite similar SFRs at the high-mass end, the efficiency of star formation in each case differs. The stellar fraction for the $\sigma_{8}$ low model is significantly boosted compared to observations, while the fiducial and $h$ high models only mildly overpredict the observed stellar fraction. The $\sigma_{8}$ high case, on the other hand, is the best match. We note that the stellar fraction is inversely proportional to $E_{\rm SN}$ for all halo masses. The weak mass dependence trend of the stellar fraction is the same compared to the unconstrained case, and in agreement with the trends of the FIREboxHR simulations (Feldmann et al., 2024) and ColdSIM (Maio & Viel, 2023) for the fiducial, $\sigma_{8}$ high and $h$ high models, while the $\sigma_{8}$ low model is an order of magnitude higher but with the same slope.
@@
While the UVLFs at $M_{\rm 1600,AB}<-18$ in the constrained case are similar for all four cosmological models, the SFR and stellar fraction show differences, and in some cases do not match observations. In such a scenario, it becomes interesting to analyze the impact on the 21-cm signal power spectrum, which is presented in the bottom panel of Figure 7 at $z=9$ for the constrained case. Unlike the unconstrained case where the differences in matter clustering contributed to differences not only in the overdensity term but also in the neutral hydrogen fraction term in Equation 2, here the changes in astrophysical parameters also contribute to the neutral hydrogen fraction and reduce the impact of matter clustering, leading to a drastic reduction in the differences between the value of $\langle x_{\rm HI}\rangle$ across the four models. Thus, the differences in the 21-cm signal power spectra of the four models are governed primarily by the overdensity term. This leads to a higher power across all wave modes for the $\sigma_{8}$ high model due to greater matter clustering. While one would expect the $h$ high model to have slightly higher power than the fiducial one due to the dependence on overdensity, the neutral-fraction term in Equation 2 reduces the contrast due to a marginally faster reionization process, so that the power in the two cases becomes similar. On the other hand, due to lower matter clustering, the $\sigma_{8}$ low model has a consistently lower power amplitude as compared to the fiducial model.
[…]
@@ 4 Discussion
By analyzing various galactic and IGM observables, we note that the interplay of astrophysical and cosmological parameters can lead to similarities as well as differences. For example, in the constrained case, while agreement with observed UVLFs leads to similar SFR for the high-mass galaxies across all models, there are significant differences in the SFR and stellar fraction for low-mass galaxies. On the other hand, while in the unconstrained case, the SFR and stellar fraction show a very similar dependence on stellar mass in all models because the parameters that regulate such quantities are the same, the UVLFs exhibit significant differences.
@@
These similarities in terms of 21-cm observables have a profound impact on inference modeling, as models with very different cosmological and astrophysical parameters may still produce 21-cm signal power spectra in agreement with current upper limits (for example from Mertens et al., 2020; Trott et al., 2020; HERA Collaboration et al., 2023; Acharya et al., 2024c), or possibly also with an eventual measurement. This is because, while the end of reionization may differ among models, at $z>7$ the extent and speed of reionization are still largely similar. This means that fixing a priori any of these parameters could exclude viable models. While introducing cosmological and astrophysical parameters as free parameters during inference modeling has been attempted in several earlier studies (see e.g. Santos & Cooray, 2006; Liu & Parsons, 2016; Kern et al., 2017; Hassan et al., 2020; Schneider et al., 2023), these have used only approximate reionization simulation codes due to the significant computational costs involved in building such a high-dimensional parameter space. Indeed, in a more physical approach, it is necessary to run a large number of dark-matter only simulations with different cosmologies, each post-processed with a large set of astrophysical parameters. The problem worsens if one would like to boost the resolution of the simulations to take into account the role of mini halos (see Haiman et al., 2001; Iliev et al., 2005), dwarf galaxies (Wu & Kravtsov, 2024), or Lyman Limit Systems (Georgiev et al., 2024; Giri et al., 2024) which likely play a significant role in the EoR. While faster semi-numerical codes and emulators can be employed, they are not ideal tools to model galaxy-scale physics and radiative transfer effects.
@@ 5 Summary
In this work, we have investigated the impact of different cosmological models on the 21-cm signal and some galactic properties using Polar, which combines $N$-body dark matter simulations run with Gadget-4 with the semi-analytic model of galaxy formation L-Galaxies, and the 1-D radiative transfer code Grizzly. We have applied the framework to four different cosmological models: a “fiducial” model, which adopts values of the cosmological parameters from Planck Collaboration et al. (2020); a “$h$ high” model, with $h=0.733$ based on results from studies of Cepheid variables in the host galaxies of 42 Type Ia supernovae (Riess et al., 2022); a $\sigma_{8}$ low” case, with $\sigma_{8}=0.702$ from an anisotropic galaxy clustering measurement analysis done by Tröster et al. (2020); and a $\sigma_{8}$ high” case, with $\sigma_{8}=0.880$ according to recent eROSITA results (Ghirardini et al., 2024). We additionally used the Fixed & Paired approach (Angulo & Pontzen, 2016) to suppress cosmic variance by boosting the effective volumes of the simulations. For all the quantities analyzed, we took an average of the fixed initial conditions case (which we refer to as the A-series) and its corresponding pair (referred to as the B-series).
We then choose astrophysical parameters in L-Galaxies such that the fiducial model matches UV luminosity functions from HST legacy fields and JWST programs at $z=10$ and 9. In particular, we focus on parameters affecting the star formation efficiency ($\alpha_{\rm SF}$), star formation efficiency during galaxy mergers ($\alpha_{\rm SF,burst}$), AGN accretion rate ($k_{\rm AGN}$), reheating of cold gas by star formation ($\epsilon_{\rm reheat}$, $V_{\rm reheat}$), and the energy released by each supernova ($E_{\rm SN}$).
To investigate the effect of adopting different values for $h$ and $\sigma_{8}$ on other galactic and IGM observables, we use the same astrophysical parameters also for the other three cosmological models. We refer to this as the “unconstrained” case, as by not changing astrophysical parameters for the different cosmological models the resulting UVLFs will not necessarily be consistent with the observed ones. We also built a “constrained” case, where we instead choose the astrophysical parameters such that the UVLFs obtained in each cosmological model are consistent with those observed at $z=10$ and 9. For this, we increased the star formation efficiency and reduced the energy released per supernova in the $\sigma_{8}$ low model to boost star formation and, in turn, its UVLF. On the other hand, we increased the energy released per supernova for the $h$ high and $\sigma_{8}$ high models to suppress star formation and mitigate the impact of increased matter clustering as compared to the fiducial model. For the radiative transfer calculations, we only took into account the stellar sources while modeling the spectral energy distributions and chose a global escape fraction of 12.5% across all redshifts for all four models.
@@
  • Global Star Formation Efficiency (SFE): As the parameters controlling star formation are the same in the unconstrained case, the global SFE estimated via the stellar fraction at all masses is similar in all models. The stellar fraction is also in agreement with observations and abundance matching estimates. However, in the constrained case, due to a higher value of the parameter regulating the star formation efficiency, the stellar fraction in the $\sigma_{8}$ low model is also higher. The other three models show minor differences, governed by the differences in the energy released per supernova.
@@
  • 21-cm signal power spectrum: In the unconstrained case, the differences between the four models are not only due to the difference in overdensity but also to those in the neutral fraction and its redshift evolution. However, in the constrained case, the neutral fraction at $z>8$ is similar in all models, as the impact of matter clustering on the neutral fraction is cancelled out by the impact of setting different values for the astrophysical parameters controlling star formation. Thus, the 21-cm signal power spectrum at higher redshifts is mostly dictated by the overdensity term, and because of this it shows smaller differences among the four models.
@@
Overall, we conclude that jointly varying cosmological and astrophysical parameters can lead to differences in some observables (e.g. low-mass SFR and stellar fraction), while others stay similar (e.g. UVLFs and 21-cm signal power spectrum with the latter remaining well within the expected observational errors despite minor differences). In particular, we note that despite significantly different galactic processes and reionization histories, the 21-cm power spectra are very similar in power across $k$-bins and in agreement with current observational upper limits (Acharya et al., 2024c; Mertens et al., 2020; Trott et al., 2020; HERA Collaboration et al., 2023). Due to this, when doing inference modeling, it is essential to consider all cosmological and astrophysical parameters as free parameters, with other observational constraints serving as priors. While a limited exploration of astrophysical parameters can be done with excursion set algorithms, a semi-analytic model provides a significantly more rigorous and physical approach to modeling galactic properties. For varying cosmological parameters though, it is necessary to run a large number of $N$-body simulations to populate the prior parameter space. However, running so many high-resolution simulations would be prohibitively expensive in terms of computational resources and time. To address this issue, in future work we intend to incorporate techniques for boosting the resolution of less costly, low-resolution simulations using either analytic or machine learning techniques. Such an implementation would also additionally allow us to build broader training sets of power spectra for the Machine Learning kernels used with Gaussian Process Regression by the LOFAR EoR Key Science Project team (as done by Acharya et al., 2024a).
@@
Lastly, to achieve a stronger constraining power on a broadened parameter space, it is essential to jointly employ a variety of galactic observables, e.g. from JWST, Euclid, and upcoming SPHEREx, as well as tomographic imaging of neutral Hydrogen in the IGM with the future SKA-Low.